The combustion of ammonia in the presence of excess oxygen yields NO2 and H2O:

4 NH3 (g) + 7 O2 (g) → 4 NO2 (g) + 6 H2O (g), The combustion of 28.8 g of ammonia consumes __________ g of oxygen.


1
Expert's answer
2021-03-04T06:34:19-0500

moles of NH3 = 28.8/17 = 1.694 moles

4 moles of NH3 requires 7 moles of O2

1 moles requires = 7/4 moles of O2

1.694 moles requires = 1.694×7/4 = 2.964 moles of O2

mass of O2 = 2.964×32 = 94.848g